About what percent of the x values from a normal distribution lie within two standard deviations


24. About what percent of the x values from a normal distribution lie within two standard deviations (left and right) of the mean of that distribution? (Enter an exact number as an integer, fraction, or decimal.) 

_______  % 

25. About what percent of x values lie between the mean and one standard deviation (one sided)? (Enter an exact number as an integer, fraction, or decimal.) 

_______ %  

26. About what percent of x values lie between the first and third standard deviations (both sides)? (Enter an exact number as an integer, fraction, or decimal.) 

_______  % 
Answer #1

Empirical rule is: 68 - 95 - 99.7 that's what we will use here.

24) 2 sd hence it's 95% by empirical rule.

25) between mean and 1 sd total is 68% in two sides, we need one sided so it's 68/2 = 34%

26) Value between 1st and 3rd sd = (99.7 - 68) = 31.7%

All these are based on empirical rule.

If any doubt please comment. Thank you.